OpenStudy (anonymous):

find the equation of the tangent line at the point specified.. y = xtanx ; x = pi/4

OpenStudy (bahrom7893):

x*(secx)^2 + tan(x), plug in x = pi/4

OpenStudy (anonymous):

take the derivative, using the product rule. replace x by \[\frac{\pi}{4}\] then use point slope formula

OpenStudy (anonymous):

yes but do i have to apply the power rule twice, once with (tanx) with x(tanx)?

OpenStudy (bahrom7893):

no pwer rules, just product rule

OpenStudy (anonymous):

sorry thats what i meant. do i have to apply it twice?

OpenStudy (bahrom7893):

no, just once. This is in form of u*v, where u = x and v = tanx (u*v)' = u*v' + v*u' = x*(Secx)^2 + (tanx)*1

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

y = xtanx ; x = pi/4 f'(x) = tanx + x ( sec^2x) = tanx + x( 1 + tan^2x) -> f'( pi/4) = tan(pi/4) + pi/4 ( 1 + tan^2pi/4) = 1 + pi/4 + pi/4 = 1 + pi/2 = 2.57 Tangent line: y - pi/4 = 2.57( x - pi/4)